Here is a question:
At an Italian restaurant there is a jug of heterogeneous oils that are composed of 20 liters of 93% oil type and 10 liters of pure neutral liquid.
Now, the 93% oils are on the bottom while the neutral liquid is less dense. How much of the oils must I take out of the bottom, so that when I shake the jug up, I get a heterogeneous solution of 64%?

Answer by Nate(3500)

First, let us find the solution's percentage of oil content:
(20(93%) + 10(0%))/(30) = ?
(20(0.93) + 10(0.00))/30 = ?
18.6/30 = 0.62 or 62%
On the contrary, you would have to add oils to get a solution of 64% ....
(18.6 + x(0.93))/(30 + x) = 0.64
18.6 + 0.93x = 19.2 + 0.64x
0.29x = 0.6
x = about 2.1 liters of 93% oils must be added